Lebron is taking Calculus for Management, a class in which four tests are given. To get an A, he must average at least a 90 on the four tests. He got scores 89, 93, and 87 on the first three tests. What is the least he could score on the last test to earn an A in the class? (Express your answer as an inequality)
Answer by Theo(13342)

he has already scored 89 + 93 + 87

the sum of this is 269.

in order to average a 90, he needs to score a total of 4 * 90 = 360 on all 4 tests.

360 - 269 = 91.

he needs to score at least a 91 to get a 90 average on all 4 tests.

269 + 91 = 360

360 / 4 = 90

the general formula is sum of scores / number of tests = average score on each test.

from this formula, you can easily derive:

sum of scores = average score on each test * number of tests

when he gets a 91 on the last test, than the 4 individual scores are 89, 93, 87, 91.

add them up and you get a total of 360 for 4 scores.

average is 360 / 4 = 90.